OpenStudy (anonymous):

One canned juice drink is 30% orange juice; another is 5% orange juice. How many liters of each should be mixed together in order to get 25L that is 8% orange juice? How do you determine how many liters of the 30% would be used and how many of the 5% would be used?

OpenStudy (anonymous):

x=liter of 30 percent solution y= liter of 5 percent solution .3x+.05y=25(.08) x+y=25 This is what I get but how do I determine the L of 30% and L 5%?

OpenStudy (anonymous):

30 L + 5* ( 25 - L) = 8 * 25 Do you follow me?

OpenStudy (anonymous):

Chlorophyll, I'm kinda of loss? can you explain plz?..

OpenStudy (anonymous):

Luis Rivera, How do you determine how many liters of the 30% would be used and how many of the 5% would be used?